The expression 2(l+W) may be used to find the perimeter of a rectangle
What are the length and the width of a rectangle if the area is 13 1/2
square unites and the lenght of one side is 1/5 the measure of the perimeter?

Answer by checkley71(8403)

.2*13.5=2.7 FOR THE LENGTH.
2*2.7+2*W=13.5
5.4+2W=13.5
2W=13.5-5.4
2W=8.1
W=4.05 WIDTH MEASUREMENT
PROOF
2*2.7+2*4.05=13.5
5.4+8.1=13.5
13.5=13.5